# Time-varying Vector Field Compression with Preserved Critical Point Trajectories

### v2

@@ III-E Face-Level Test
Consider a triangular face $f$ of the space–time tetrahedral mesh. Let the three incident vertices carry vector values $\{\mathbf{a},\mathbf{b},\mathbf{c}\}\subset\mathbb{R}^{2}$. Because $\mathbf{V}$ is piecewise-linear, according to (1), the restriction of $\mathbf{V}$ to $f$ is the convex hull of $\{\mathbf{a},\mathbf{b},\mathbf{c}\}$ in $\mathbb{R}^{2}$ , which can be written as:
@@ III-F Simulation of Simplicity (SoS)
III-G Derivation of error bounds for critical point preservation
@@ IV-A Research Problem and Goal
$\|\hat{\mathbf{D}}-\mathbf{D}\|_{\infty}\leq\varepsilon,$
@@
For every space–time triangular face $f=\{i,j,k\}$ (including both time-slice faces and cross-time-slab faces), define the face-level critical point predicate
$P_{f}(\mathbf{d}_{i},\mathbf{d}_{j},\mathbf{d}_{k}):=\mathbb{I}\!\big(\ \mathbf{0}\in\mathrm{conv}\{\mathbf{d}_{i},\mathbf{d}_{j},\mathbf{d}_{k}\}\ \big),$
@@
where $\mathcal{T}(\cdot)$ denotes the track set, i.e., the collection of space–time trajectories obtained by connecting critical points across consecutive timesteps through their face crossings in the space–time mesh.
@@ IV-B Guarantees
If the critical point predicate on every triangular face of the space–time mesh is invariant under compression, then for every time slice $t_{k}$, the critical point set on $D\times\{t_{k}\}$ is unchanged (same cardinality and locations via piecewise-linear barycentric mapping).
Sketch.
@@
Sketch.
Within a tetrahedron, the two boundary intersections determine a unique segment; preserving which faces are crossed preserves which segment exists. Across tets, two segments meet if and only if they share the same crossing point on a shared face. Because the face predicate and the barycentric crossing on that face are preserved, the segment gluing is unchanged, yielding an isomorphic track graph. ∎
According to Theorem 1 and Theorem 2, we can conclude that preserving all per-time critical points as well as all cross-time-slab critical points constitutes a sufficient condition for ensuring the preservation of critical point trajectories.
@@ V Design Overview
This section presents an overview of our framework and describes how the proposed method preserves critical-point trajectories during lossy compression.
@@
Algorithm 3 Critical Point Trajectories Preserving Lossy Compression for 2D Temporal Vector Fields
@@
Algorithm 4 Derive the minimum error bound for vertex $v$
@@
Figure 5: An illustrative example of Critical Point Trajectories Preserving Lossy Compression, as described in Algorithm 3 (lines 10–17).
@@ V-A Framework Overview
As illustrated in Figure 4, our compression pipeline is built on the prior work cpSZ-sos [xia2024preserving], with several key enhancements and modifications. The input data are fed into the space-time discretization module, as discussed in Section III-C, which enables our algorithm to handle time-varying 2D data by constructing a consistent spatio-temporal mesh representation. The framework records two components: (1) the critical-point-preserving error bounds associated with each vertex, and (2) the residuals between the original data and their predicted values. These two components are quantized separately and encoded using Huffman coding, followed by an additional lossless compression stage to further improve the overall compression ratio.
For the critical-point-preserving error bound component, the key difference between our method and prior work lies in how temporal information is handled. By adopting a fixed tetrahedralization strategy, we explicitly consider all triangular faces formed between adjacent time slabs, which allows us to track how critical points enter and exit cells across consecutive time steps. As a result, our method preserves critical point trajectories in time-varying data, extending previous approaches that were limited to static settings.
@@
We present our critical point trajectories preserving lossy compression in Algorithm 3. We first convert the input data into a fixed-point representation so that the SoS method can be applied. The data are then tetrahedralized and partitioned according to a predefined strategy. Next, a critical point test is performed on all triangular faces of each tetrahedron, and the results are recorded. For each vertex, we derive an error bound that guarantees the preservation of critical points based on its incident triangular faces (up to 36 faces), and quantize the vertex value accordingly. Subsequently, prediction and quantization are applied to the vector data at this vertex. Then, we compute the decompressed data and replace the original value with it on the fly, ensuring that subsequent processing operates on previously decompressed results. In this way, the compression process is performed sequentially, and the processing of each data point explicitly depends on the decompressed values of preceding points. Finally, all quantized error bounds, quantized vector data, and unpredictable data that must be stored losslessly are packaged together and passed to a lossless compressor for final encoding.
@@ V-B Time Complexity Analysis
Denote $N$ as the total number of data points, and we analyze the complexity of our compression algorithm as follows. Our algorithm consists of three stages. In the first stage, we perform a robust critical point test on all triangles incident to each vertex. Each test has a time complexity of $O(m)$, where $m$ denotes the dimensionality ($m=2$ for the 2D datasets). According to our prism decomposition strategy, each vertex is incident to at most 36 triangles. Therefore, the complexity of this stage is $O(\alpha mN)$, where $\alpha=36$ denotes the maximal number of triangular faces associated with a data point. Since both $\alpha$ and $m$ are constants, this stage has an overall time complexity of $O(N)$. As a variant of the SZ compressor family, prior work shows this stage has complexity $O(N)$ [sz16]. The third stage derives error bounds for preserving critical point trajectories. Similar to the first stage, it requires per-vertex analysis over incident triangles and therefore has complexity $O(\alpha N)=O(N)$. Overall, the total time complexity of our algorithm is $O(N)$.
@@
We focus on compressing 2D temporal vector field data $\mathbf{v}(x,y,t)=(u,v)$. Numerical simulation data often exhibit strong local physical heterogeneity. In different subdomains, the dominant physical mechanisms such as advection, convection, nonlinear transport, and diffusion can vary significantly. Most state-of-the-art scientific lossy compressors [sz17, sz16, sz18, liang2018efficient] adopt the first-order Lorenzo predictor as their prediction model. We refer readers to [ibarria2003out] for further details. This model performs well in locally smooth or diffusion-dominated regions due to its trilinear reproduction property, but it produces large residuals in regions with significant translational or transport structures (for example, vortex-core migration or stripe advection), which reduces compression efficiency.
To address this limitation, we propose a block-wise adaptive Mixture of Predictors with two candidates: First order 3D-Lorenzo(3DL) and Semi-Lagrangian (SL) predictor tailored for advection-dominated regions. We partition each frame at time $t$ into non-overlapping blocks $B$ of size $B_{x}\times B_{y}$. For each spatial block, we evaluate both predictors with a light-weight, context-preserving look-ahead, then select the mode that minimizes the estimated rate; this reduces residual entropy and improves overall compression ratio while strictly preserving the user-prescribed error bound.
@@ VI-A Semi-Lagrangian Predictor
Let spatial steps be $\Delta x,\Delta y$, time step $\Delta t$(these parameters are included in the data’s metadata or descriptive information), and denote the fixed-point scaling by $S$ (the data are stored as $S$-scaled integers). Define per-axis Courant–Friedrichs–Lewy (CFL)-like factors
$\mathrm{CFL}_{x}=\frac{\Delta t}{\Delta x}\cdot\frac{1}{S}\qquad\mathrm{CFL}_{y}=\frac{\Delta t}{\Delta y}\cdot\frac{1}{S}$
@@
$\big(i^{\star},j^{\star}\big)$ $=\Big(i-v_{t-1}(i,j)\,\mathrm{CFL}_{y},\;j-u_{t-1}(i,j)\,\mathrm{CFL}_{x}\Big)$ (5)
@@
$(i_{\frac{1}{2}},j_{\frac{1}{2}})=\Big(i-\tfrac{1}{2}v_{t-1}(i,j)\,\mathrm{CFL}_{y},\;j-\tfrac{1}{2}u_{t-1}(i,j)\,\mathrm{CFL}_{x}\Big),$ (7)
@@
$(i^{\star},j^{\star})=\Big(i-v_{t-1}(i_{\frac{1}{2}},j_{\frac{1}{2}})\,\mathrm{CFL}_{y},\;j-u_{t-1}(i_{\frac{1}{2}},j_{\frac{1}{2}})\,\mathrm{CFL}_{x}\Big)$ (8)
@@
$(i_{s+1},j_{s+1})$ $=(i_{s},j_{s})-\Big(v_{t-1}(i_{s},j_{s})\,\tfrac{\mathrm{CFL}_{y}}{N},\;u_{t-1}(i_{s},j_{s})\,\tfrac{\mathrm{CFL}_{x}}{N}\Big)$
@@
After we obtain $(i^{\star},j^{\star})$, use equation (6) to predict the $u$ or $v$ component.
@@
Figure 6 provides a schematic illustration of two prediction methods we proposed. The red point indicates the target location to be predicted. We first compute the coordinates of all intermediate steps (yellow points) using (7) or (9). Then, for each intermediate step, we perform bilinear interpolation of the components $u$ and $v$ at time t1t-1 from its four surrounding points to obtain the local velocity $(u,v)$ at the current intermediate position. This process is repeated iteratively until the final position $(i^{\star},j^{\star})$ is reached, where we again perform bilinear interpolation of the four neighboring points at time t1t-1 to obtain the predicted value.
@@ VI-B Data Sampling and Estimation
$H_{0}(h_{p};n_{p})\;=\;-\sum_{k}\frac{h_{p}[k]}{n_{p}}\,\log_{2}\!\Big(\frac{h_{p}[k]}{n_{p}}\Big),\qquad n_{p}>0$ (10)
@@
$\widehat{R}_{p}(B)\;=\;H_{0}(h_{p};n_{p})\;+\;\lambda\,\widehat{\xi}_{p}\;+\;R_{\mathrm{meta}},\;\;\widehat{\xi}_{p}\;=\;\frac{\mathcal{L}_{p}}{\,n_{p}+\mathcal{L}_{p}\,}$
[…]
@@ VII-A Experimental Setup and Datasets
We conducted our experiments using five scientific datasets listed in Table I, where #CP denotes the total number of critical points located at time $t$, and within a time slab. All datasets [Jung93, gerrisflowsolver, Guenther17, Jakob20, Basilisk] are stored as single-precision floating-point values with the following detailed information.
@@
  • Fluid Simulation Ensemble(FS): An ensemble simulation of 8000 unsteady two-dimensional flows, with turbulence becoming more prominent in the later snapshots. Without loss of generality, we randomly pick the $6666^{th}$ flow dataset for evaluation.
@@ VII-B Baselines
To the best of our knowledge, apart from lossless compression, no existing lossy compressor is capable of preserving the critical point trajectories in temporal 2D vector field data. Since lossless compressors such as Gzip, Zstd, and FPZIP naturally guarantee full topological fidelity, we use them as upperbound baselines for comparison. In addition, we evaluate three representative lossy compressors: ZFP, SZ3, and cpSZ-SoS to investigate how lossy techniques perform in terms of critical point trajectories preservation. All evaluated compressors operate in $O(N)$.
@@ VII-C Metrics
  • Compression Ratio (CR)
  • Peak Signal-to-Noise Ratio (PSNR): Evaluates overall field fidelity in the value domain, defined as $\text{PSNR}=20\cdot\log_{10}(\text{data range})-10\cdot\log_{10}(\text{MSE})$.
  • False Cases of Critical Points: This metric quantifies the incorrect detection or correspondence of critical points, including both false positives (spurious critical points) and false negatives (missing critical points). It is measured at two temporal scales: (a) Spatial level at time $t$ ($FC_{t}$). (b) Space–time level at time slab $t_{k},t_{k+1}]$ ($FC_{s}$).
  • Number of Trajectories: The number of critical point trajectories formed by connecting critical points detected at each time step $t$ through their correspondences across consecutive time slabs.
Figure 8: Rate–distortion curves of the three methods on five datasets under varying error bounds. The x-axis denotes the bit rate (data type size in bits divided by the compression ratio), and the y-axis denotes PSNR. CFVKV, FS, and HCBA use [5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1]; SCF uses [1e-2, 5e-2, 1e-1, 0.5, 1, 5]; and DT uses [1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2]. At the same PSNR, a lower bit rate indicates better compression performance.
@@
VII-D Rate-distortion
Figure 8 presents the rate–distortion curves of our three methods across five datasets, explicitly illustrating the sensitivity to different error bounds. For each dataset, the error bounds are selected relative to the data range. We observe that, under the same PSNR, the SL method performs exceptionally well in terms of compression ratio on the CFVKV and FS datasets when the error bounds are relatively large, while the 3DL method achieves better compression ratios on the SCF and FS datasets. In contrast, the MoP method effectively combines the strengths of both approaches, achieving a compression ratio that either surpasses or closely matches the best of the other two methods in most cases.

### v3

@@ III-E Face-Level Test
Consider a triangular face $f$ of the space-time tetrahedral mesh. Let the three incident vertices carry vector values $\{\mathbf{a},\mathbf{b},\mathbf{c}\}\subset\mathbb{R}^{2}$. Because $\mathbf{V}$ is piecewise-linear, according to (1), the restriction of $\mathbf{V}$ to $f$ is the convex hull of $\{\mathbf{a},\mathbf{b},\mathbf{c}\}$ in $\mathbb{R}^{2}$ , which can be written as:
@@ III-F Simulation of Simplicity
III-G Derivation of Error Bounds for Critical-Point Preservation
@@ IV-A Research Problem and Goal
$\|\hat{\mathbf{D}}-\mathbf{D}\|_{\infty}\leq\varepsilon;$
@@
For each space-time triangular face $f=\{i,j,k\}$ (including both time-slice faces and cross-time-slab faces), we define a face-level critical-point predicate (hereafter referred to as the face predicate)
$P_{f}(\mathbf{d}_{i},\mathbf{d}_{j},\mathbf{d}_{k}):=\mathbb{I}\big(\mathbf{0}\in\mathrm{conv}\{\mathbf{d}_{i},\mathbf{d}_{j},\mathbf{d}_{k}\}\big).$
@@
where $\mathcal{T}(\cdot)$ denotes the track set, i.e., the collection of space-time trajectories obtained by connecting critical points across consecutive timesteps through their face crossings in the space-time mesh.
@@ IV-B Guarantees
If the face-level critical-point predicate on every triangular face of the space-time mesh is invariant under compression, then for every time slice $t_{k}$, the critical point set on $D\times\{t_{k}\}$ is unchanged (same cardinality and locations via piecewise-linear barycentric mapping).
Proof Sketch.
@@
Proof Sketch.
Within a tetrahedron, the two boundary intersections uniquely determine a segment; thus, preserving which faces are intersected ensures that the same segment is retained. Across tetrahedra, two segments connect if and only if they share the same intersection point on a common face. Since both the face predicate and the corresponding barycentric intersection on that face are preserved, the segment connectivity remains unchanged, yielding an isomorphic track graph. ∎
According to Theorem 1 and Theorem 2, we can conclude that preserving all per-time critical points as well as all cross-time-slab critical points constitutes a sufficient condition for ensuring the preservation of critical-point trajectories.
@@ V Design Overview
This section provides an overview of our framework and outlines how the proposed method preserves critical-point trajectories during lossy compression.
@@
Algorithm 3 Lossy compression for 2D time-varying vector fields with critical-point trajectory preservation
@@
Algorithm 4 Derive the minimum error bound for a vertex
@@
Figure 5: An illustrative example of critical-point-trajectory-preserving lossy compression, as described in Algorithm 3 (lines 10–17).
@@ V-A Framework Overview
As illustrated in Figure 4, our compression pipeline is built on the prior work cpSZ-SoS [xia2024preserving], with several key enhancements and modifications. The input data are fed into the space-time discretization module, as discussed in Section III-C, which enables our algorithm to handle time-varying 2D data by constructing a consistent spatiotemporal mesh representation. The framework records two components: (1) the critical-point-preserving error bounds associated with each vertex, and (2) the residuals between the original data and their predicted values. These two components are quantized separately and encoded using Huffman coding, followed by an additional lossless compression stage to further improve the overall compression ratio.
For the critical-point-preserving error bound component, the key difference between our method and prior work lies in how temporal information is handled. By adopting a fixed tetrahedralization strategy, we explicitly consider all triangular faces formed between adjacent time slabs, which allows us to track how critical points enter and exit cells across consecutive time steps. As a result, our method preserves critical-point trajectories in time-varying data, extending previous approaches that were limited to static settings.
@@
We present our critical-point-trajectory-preserving lossy compression in Algorithm 3. We first convert the input data into a fixed-point representation so that the SoS method can be applied. The data are then tetrahedralized and partitioned according to a predefined strategy. Next, a critical point test is performed on all triangular faces of each tetrahedron, and the results are recorded. For each vertex, we derive an error bound that guarantees the preservation of critical points based on its incident triangular faces (up to 36 faces), and quantize the vertex value accordingly. Subsequently, prediction and quantization are applied to the vector data at this vertex. Then, we compute the decompressed data and replace the original value with it on the fly, ensuring that subsequent processing operates on previously decompressed results. In this way, the compression process is performed sequentially, and the processing of each data point explicitly depends on the decompressed values of preceding points. Finally, all quantized error bounds, quantized vector data, and unpredictable data that must be stored losslessly are packaged together and passed to a lossless compressor for final encoding.
@@ V-B Time Complexity Analysis
Denote $N$ as the total number of data points, and we analyze the complexity of our compression algorithm as follows. Our algorithm consists of three stages. In the first stage, we perform a robust critical point test on all triangles incident to each vertex. Each test has a time complexity of $O(m)$, where $m$ denotes the dimensionality ($m=2$ for the 2D datasets). According to our prism decomposition strategy, each vertex is incident to at most 36 triangles. Therefore, the complexity of this stage is $O(\alpha mN)$, where $\alpha=36$ denotes the maximal number of triangular faces associated with a data point. Since both $\alpha$ and $m$ are constants, this stage has an overall time complexity of $O(N)$. As a variant of the SZ compressor family, prior work shows this stage has complexity $O(N)$ [sz16]. The third stage derives error bounds for preserving critical-point trajectories. Similar to the first stage, it requires per-vertex analysis over incident triangles and therefore has complexity $O(\alpha N)=O(N)$. Overall, the total time complexity of our algorithm is $O(N)$.
@@ VI Block-Wise Adaptive Mixture of Predictors
We focus on compressing 2D time-varying vector field data $\mathbf{v}(x,y,t)=(u,v)$. Numerical simulation data often exhibit strong local physical heterogeneity. In different subdomains, the dominant physical mechanisms such as advection, convection, nonlinear transport, and diffusion can vary significantly. Most state-of-the-art scientific lossy compressors [sz17, sz16, sz18, liang2018efficient] adopt the first-order Lorenzo predictor as their prediction model. We refer readers to [ibarria2003out] for further details. This model performs well in locally smooth or diffusion-dominated regions due to its trilinear reproduction property, but it produces large residuals in regions with significant translational or transport structures (for example, vortex-core migration or stripe advection), which reduces compression efficiency.
To address this limitation, we propose a block-wise adaptive mixture of predictors (MoP) with two candidates: first-order 3D-Lorenzo (3DL) predictor and semi-Lagrangian (SL) predictor tailored for advection-dominated regions. We partition each frame at time $t$ into non-overlapping blocks $B$ of size $B_{x}\times B_{y}$. For each spatial block, we evaluate both predictors with a light-weight, context-preserving look-ahead, then select the mode that minimizes the estimated rate; this reduces residual entropy and improves overall compression ratio while strictly preserving the user-prescribed error bound.
@@ VI-A Semi-Lagrangian Predictor
Let the spatial steps be $\Delta x$ and $\Delta y$, and the time step be $\Delta t$; these parameters are provided in the data metadata or descriptive information. Let $S$ denote the fixed-point scaling factor, such that the data are stored as integers scaled by $S$. Define per-axis Courant–Friedrichs–Lewy (CFL)-like factors,
$\mathrm{CFL}_{x}=\frac{\Delta t}{\Delta x}\cdot\frac{1}{S},\qquad\mathrm{CFL}_{y}=\frac{\Delta t}{\Delta y}\cdot\frac{1}{S}.$
@@
$\big(i^{\star},j^{\star}\big)$ $=\Big(i-u_{t-1}(i,j)\,\mathrm{CFL}_{x},\;j-v_{t-1}(i,j)\,\mathrm{CFL}_{y}\Big)$ (5)
@@
$(i_{\frac{1}{2}},j_{\frac{1}{2}})=\Big(i-\tfrac{1}{2}u_{t-1}(i,j)\,\mathrm{CFL}_{x},\;j-\tfrac{1}{2}v_{t-1}(i,j)\,\mathrm{CFL}_{y}\Big),$ (7)
@@
$(i^{\star},j^{\star})=\Big(i-u_{t-1}(i_{\frac{1}{2}},j_{\frac{1}{2}})\,\mathrm{CFL}_{x},\;j-v_{t-1}(i_{\frac{1}{2}},j_{\frac{1}{2}})\,\mathrm{CFL}_{y}\Big)$ (8)
@@
$(i_{s+1},j_{s+1})$ $=(i_{s},j_{s})-\Big(u_{t-1}(i_{s},j_{s})\,\tfrac{\mathrm{CFL}_{x}}{N},\;v_{t-1}(i_{s},j_{s})\,\tfrac{\mathrm{CFL}_{y}}{N}\Big)$
@@
Figure 6 provides a schematic illustration of the two proposed prediction methods. The red point indicates the target location to be predicted. We first compute the coordinates of all intermediate steps (yellow points) using (7) or (9). Then, for each intermediate step, we perform bilinear interpolation of the components $u$ and $v$ at time t1t-1 from its four surrounding points to obtain the local velocity $(u,v)$ at the current intermediate position. This process is repeated iteratively until the final position $(i^{\star},j^{\star})$ is reached, where we again perform bilinear interpolation of the four neighboring points at time t1t-1 to obtain the predicted value.
@@ VI-B Data Sampling and Estimation
$H_{0}(h_{p};n_{p})\;=\;-\sum_{k}\frac{h_{p}[k]}{n_{p}}\,\log_{2}\!\Big(\frac{h_{p}[k]}{n_{p}}\Big),\qquad n_{p}>0.$ (10)
@@
$\widehat{R}_{p}(B)\;=\;H_{0}(h_{p};n_{p})\;+\;\lambda\,\widehat{\xi}_{p}\;+\;R_{\mathrm{meta}},\;\;\widehat{\xi}_{p}\;=\;\frac{\mathcal{L}_{p}}{\,n_{p}+\mathcal{L}_{p}\,}.$
Here, $\lambda$ denotes a fixed penalty measured in bits. We set it to half the data type size, assuming that subsequent lossless compression achieves an average compression ratio of $2\times$ [lindstrom2017error, son2014data]. Since only two modes are considered, the metadata overhead is one bit per block. We select the mode with the lower estimated score, subject to a relative-improvement criterion. Specifically, if the improvement of $\widehat{R}_{p}^{SL}(B)$ exceeds a switching threshold (set to $0.3\%$), we adopt the SL mode for that block; otherwise, we revert to 3DL to avoid mode thrashing caused by noisy estimates.
@@ VII-A Experimental Setup and Datasets
We conduct our experiments using five scientific datasets listed in Table I, where #CP denotes the total number of critical points located at time $t$, and within a time slab. All datasets [Jung93, gerrisflowsolver, Guenther17, Jakob20, Basilisk] are stored as single-precision floating-point values with the following detailed information.
@@
  • Fluid Simulation Ensemble (FS): An ensemble simulation of 8000 unsteady 2D flows, with turbulence becoming more prominent in the later snapshots. Without loss of generality, we randomly pick the $6666^{th}$ flow dataset for evaluation.
@@ VII-B Baselines
To the best of our knowledge, apart from lossless compression, no existing lossy compressor is capable of preserving the critical-point trajectories in time-varying 2D vector field data. Since lossless compressors such as GZIP, ZSTD, and FPZIP inherently guarantee full topological fidelity, we use them as upper-bound baselines for comparison. In addition, we evaluate three representative lossy compressors, that is, ZFP, SZ3, and cpSZ-SoS, to assess how lossy techniques perform in preserving critical-point trajectories. All evaluated compressors have $O(N)$ time complexity.
@@ VII-C Metrics
  • Compression Ratio (CR): CR measures the data reduction capability, as defined in Section IV.
  • Peak Signal-to-Noise Ratio (PSNR): PSNR evaluates overall field fidelity in the value domain, which is defined as $\text{PSNR}=20\cdot\log_{10}(\text{data range})-10\cdot\log_{10}(\text{MSE})$.
  • False Cases of Critical Points: This metric quantifies the incorrect detection or correspondence of critical points, including both false positives (spurious critical points) and false negatives (missing critical points). It is measured at two temporal scales: (a) spatial level at time $t$ ($FC_{t}$); (b) space-time level at time slab $t_{k},t_{k+1}]$ ($FC_{s}$).
  • Number of Trajectories: The number of critical-point trajectories formed by connecting critical points detected at each time step $t$ through their correspondences across consecutive time slabs.
Figure 8: Rate-distortion curves of the three methods on five datasets under varying error bounds. The x-axis denotes the bit rate (data type size in bits divided by the compression ratio), and the y-axis denotes PSNR. CFVKV, FS, and HCBA use [5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1]; SCF uses [1e-2, 5e-2, 1e-1, 0.5, 1, 5]; and DT uses [1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2]. At the same PSNR, a lower bit rate indicates better compression performance.
@@
VII-D Ablation Study on Rate-Distortion Performance
Figure 8 presents the rate-distortion curves of our three methods across five datasets, explicitly illustrating the sensitivity to different error bounds. For each dataset, the error bounds are selected relative to the data range. We observe that, under the same PSNR, the SL method performs exceptionally well in terms of compression ratio on the CFVKV and FS datasets when the error bounds are relatively large, while the 3DL method achieves better compression ratios on the SCF and FS datasets. In contrast, the MoP method effectively combines the strengths of both approaches, achieving a compression ratio that either surpasses or closely matches the best of the other two methods in most cases.